Why is a water tower generally placed on top of a hill? If the hill is 50 m high, can it supply water to the top floor of a 100 m tall office building? If not, how does water get to that floor?


1
Expert's answer
2021-05-07T08:12:57-0400

Blaise Pascal in the 17th century proved that the pressure exerted on a water molecule is transmitted in a liquid in all directions.

As a result, Simon Stevin found that the liquid level is the same in containers containing a homogeneous liquid and connected well below the top of the liquid. This is how the law on communicating vessels was formulated: In communicating vessels filled with a homogeneous liquid, the pressure at all points of the liquid located in the same horizontal plane is the same regardless of the shape of the vessels, and the surfaces of the liquid in the communicating vessels (open at the top) are set at the same level.



The water tower provides water to reservoirs located at a height not exceeding the height of the tower. If the water tower is located on a hill, the maximum water delivery height can be increased. Under the influence of gravity at the same atmospheric pressure, water from the water tower will flow into the reservoirs located in the office building up to a height of 50 m.



Thus, a water tower at a height of 50 m is capable of supplying water to an office building up to a floor corresponding to 50 m.

To ensure floors above 50 m, additional pressure must be created in the water tower.


"P-P_{at}=\rho g h=997\frac{kg}{m^3} 9.81 \frac{m}{s^2}50m\approxeq489kPa"

In practice, as a rule, a pumping station is installed between the water tower and the office building, which creates additional pressure for the water entering the office building in order to raise the water level to a height of 100 m.

Another way to provide an office building with water is to increase the height of the water tower, i.e. find a hill with a height of 100m.